### v1

# Privacy-Preserving Retrieval Augmented Generation with Differential Privacy

Abstract
With the recent remarkable advancement of large language models (LLMs), there has been a growing interest in utilizing them in the domains with highly sensitive data that lies outside their training data. For this purpose, retrieval augmented generation (RAG) is particularly effective—it assists LLMs by directly providing relevant information from the external knowledge sources. However, without extra privacy safeguards, RAG outputs risk leaking sensitive information from the external data source. In this work, we explore RAG under differential privacy (DP), a formal guarantee of data privacy. The main challenge with differentially private RAG is how to generate long accurate answers within a moderate privacy budget. We address this by proposing an algorithm that smartly spends privacy budget only for the tokens that require the sensitive information and uses the non-private LLM for other tokens. Our extensive empirical evaluations reveal that our algorithm outperforms the non-RAG baseline under a reasonable privacy budget of $\epsilon\approx 10$ across different models and datasets.
1 Introduction
Large language models (LLMs) have shown a great deal of promise in a variety of applications. In particular, a major application of LLMs is in question-answering. The practical adoption of these systems often involves domains whose data is highly sensitive. For instance, healthcare institutions might want to utilize their internal medical records to provide precise medical information and personal feedback, while legal firms can leverage their case archives to assist clients with legal research and documentation. One way to achieve such domain-specific question-answering is through retrieval augmented generation (RAG) [5, 23, 16]. Here, we have a set of domain-specific documents; while answering a question, RAG retrieves a list of relevant documents and inputs them to LLMs as the context. However, even though this is effective for question-answering, RAG on a sensitive corpus can still leak private information about individual documents in this corpus [39, 30, 4, 29]. This is particularly problematic when end users are outside the data-holding entity, such as patients interacting with a healthcare institution’s RAG system.
Our goal in this paper is to prevent the information leakage of the sensitive external corpus by designing a privacy-preserving RAG system. For this purpose, we use differential privacy (DP) [9, 10] as a notion of privacy. Differential privacy guarantees privacy by ensuring that the participation of a single person’s data does not make much difference to the probability of any output. In our system, we assume that each RAG document comes from a single individual, and our goal is to ensure differential privacy on the eventual answer of the LLM.
There are two aspects of challenges with designing an effective RAG algorithm under DP. The first is how to fit differential privacy into the RAG framework, and the second is how to manage the privacy-utility tradeoffs. We address the first challenge by proposing an algorithm, DPVoteRAG, based on the sample-and-aggregate framework in DP [26]. Our algorithm prepares multiple LLM instances, or voters, feeds disjoint partitions of the sensitive corpus to them, and produces output tokens one by one each through the majority vote of the voters’ token outputs. Note however that LLMs often output many tokens in response to a question. This is detrimental to privacy—the composition property of differential privacy states that multiple calculations based on the same dataset leads to higher privacy degradation. To resolve this challenge, we design another algorithm, DPSparseVoteRAG, that spends a privacy budget only when we need to. More specifically, we take advantage of the fact in RAG that LLMs require sensitive corpus only when generating tokens related to the knowledge. When not, outputs from LLMs without any context suffice. We formalize this idea with the sparse vector technique in DP [11, 12]—when voters agree with the non-private output of the LLM without contexts, we will simply output the non-private one without incurring a privacy budget. Consequently, our algorithm successfully generates sufficiently long, accurate responses under a reasonable privacy budget.
We conduct extensive experiments with a series of LLMs on multiple benchmarking datasets to evaluate our algorithms. The results demonstrate that our algorithms are able to enhance the LLMs by RAG while ensuring privacy for the external corpus. We further show that DPSparseVoteRAG improves DPVoteRAG by only spending a privacy budget when necessary and enabling us to generate longer answers within a reasonable privacy budget of $\epsilon\approx 10$.
2 Preliminaries & Problem Setting
2.1 Retrieval Augmented Generation with Large Language Model
Retrieval augmented generation (RAG) is a technique to improve the performance of large language models (LLMs) on knowledge-intensive tasks by providing external knowledge. Given a question prompt, a retriever finds relevant documents from the external data source. Then, the relevant documents are added to the prompt as the contexts. An LLM (or generator) takes the augmented prompt as input and outputs the answer.
More formally, let $x\in\bigcup_{t=1}^{\infty}\mathcal{V}^{t}$ be a prompt, where $\mathcal{V}$ is some vocabulary. We further let $D$ be a dataset of documents as an external corpus with size $|D|=n$. A retriever $R$ finds a subset of $D$, $D_{x}\subset D$ that are relevant to $x$, i.e., $D_{x}=R(x,D)$. Finally, an LLM generates an answer $y=\mathrm{LLM}(x,D_{x})\in\bigcup_{t=1}^{\infty}\mathcal{V}^{t}$. The answer generation can further be decomposed into next-token generation. In particular, for each $t$, the $t$-th token $y_{t}\in\mathcal{V}$ is generated by $\mathrm{LLM}_{t}$, which takes $x$, $D_{x}$, and previously generated tokens $y_{<t}$ as inputs: $y_{t}=\mathrm{LLM}_{t}(x,D_{x},y_{<t})$.
2.2 Differential Privacy
Differential privacy is a strong cryptographically motivated definition of individual-level privacy. It guarantees that the participation of a single individual in a dataset does not change the probability of any outcome by much. In particular, suppose we have two datasets $D$ and $D^{\prime}$, each consisting of private data from $n$ individuals. We say that $D$ and $D^{\prime}$ are neighboring if they differ in a single individual’s private data. A randomized algorithm satisfies differential privacy if the output distributions on any pair of neighboring datasets are close. The formal definition is given as follows.
Definition 1 ($(\epsilon,\delta)$-Differential Privacy [9]).
A randomized algorithm $M$ satisfies $(\epsilon,\delta)$-differential privacy if for any two neighboring datasets $D,D^{\prime}$ and for any $S\subseteq\mathrm{range}(M)$,
$\Pr[M(D)\in S]\leq\exp(\epsilon)\Pr[M(D^{\prime})\in S]+\delta.$
One of the key properties of differential privacy is composition—sequential runs of differentially private algorithms also satisfy differential privacy. The composition property quantitatively captures the intuition that the more we release the information about the sensitive data, the worse the privacy guarantee becomes. More specifically, suppose $M_{1},\ldots,M_{T}$ are $(\epsilon_{0},\delta_{0})$-differentially private algorithms, which can be chosen adaptively based on previous outputs. Sequential composition theorem [9] states that the composed sequence of such algorithms guarantee $(T\epsilon_{0},T\delta_{0})$-differential privacy. Furthermore, advanced composition theorem [13, 19] states that the total privacy guarantee has $\epsilon=\mathcal{O}(\sqrt{T}\epsilon_{0})$.
2.2.1 Sparse Vector Technique
The sparse vector technique [11, 12] has originally emerged as the alternative of the composition in DP when we have so large number of numerical queries that the composition theorem of DP cannot provide a reasonable privacy guarantee but we are only interested in answers above some threshold. In such a case, the sparse vector technique algorithm, Sparse, reports whether each (noisy) query answer exceeds the threshold. It is shown that the privacy guarantee degrades by the number of queries above the threshold, instead of the total number of queries. Therefore, we save privacy budget by much when we expect only a few queries will be above the threshold. One can also report the (noisy) query outputs above the threshold by compromising privacy budget—the variant is called NumericSparse. The privacy guarantee follows by the composition of Sparse and differentially private query answering.
2.2.2 Differentially Private Generation via Sample-and-Aggregate
There has been a body of work on generating a token sequence by LLM with differential privacy. One common way to is to borrow the idea of the sample-and-aggregate framework in DP [26]. To generate a single token, a set of LLMs, each depending on a disjoint subset of the sensitive dataset $D$, generates a token or a probability distribution of a token respectively. When a set of tokens is generated, it forms an aggregate histogram of tokens, which is then carefully randomized with noise and only the most frequent token in the noisy histogram is published as the final output. When a set of probability distribution is generated, it forms a new mixture of probability distribution satisfying differential privacy. The final output token is then sampled from the resulting mixture distribution. The repetition of this process along with the composition theorem of DP yields the differentially private token sequence generation.
2.3 Problem Setting
Figure 1: Overview of our problem setting. Note that the LLM in RAG is trained outside the privacy barrier by DP.
Our goal is to generate an LLM answer to a prompt $x$ with retrieved external knowledge, $D_{x}=R(x,D)$, from a sensitive data source $D$ with differential privacy guarantee. More specifically, let the sensitive data source $D$ be a collection of individuals’ records—one record corresponds to one individual’s sensitive data [^1] It is straightforward to extend the setting to where multiple records correspond to one individual’s data by modifying the granularity of neighboring datasets in DP possibly with overhead in privacy-utility tradeoff.. We consider a realistic adversary who does not have direct access to the data source $D$ but have a capability of querying to the RAG system with any prompt $x$. We further assume that the LLM used in the RAG system is a copy of publicly available LLMs and is already pre-trained (and fine-tuned) with data disjoint from the sensitive data source $D$. That is, having access to the LLM parameters and/or pre-training (and fine-tuning) data does not provide any information on the sensitive data source $D$. To this end, we aim to formally guarantee that given any question $x$, a randomized LLM generation algorithm with RAG, $\mathrm{LLM}_{\mathrm{priv}}(x,R(x,D))$ satisfies $(\epsilon,\delta)$-differential privacy w.r.t the external knowledge data source $D$. We present the figure for this problem setting in Figure 1.
3 Differentially Private Retrieval Augmented Generation with Sparse Vector Technique
Figure 2: Overview of DPSparseVoteRAG when generating $t$-th token. It takes a question $x$, an interim answer up to $t-1$-th tokens, $y_{<t}$, and sensitive external knowledge $D$ as inputs and outputs a new token $y_{t}$ or $y_{t}^{\mathrm{non-RAG}}$. Components enclosed in red indicate that operations are done in a privacy-preserving manner. DPVoteRAG works the same until computing the output token histogram $h_{t}$, but it immediately finds the most frequent element in $h_{t}$ afterwards. Note that each retriever output only depends on $x$ and the corresponding subset of $D$; thus, the algorithm needs not to repeat the retrieval process for every token generation. We add the retrieval procedure in the figure for completeness.
Our differentially private RAG algorithm consists of two main components—DP voting for the single-token generation and efficient privacy budget spending by leveraging the sparse vector technique combined with the utilization of LLMs without any relevant documents provided. These two components enable us to generate answers that incorporate external knowledge while guaranteeing reasonable level of differential privacy. We start from our algorithm with the first component alone, and then extend it to include the second component. The graphical overview of our algorithm is presented in Figure 2.
3.1 DPVoteRAG: Differentially Private Voting Algorithm for RAG
Input: Prompt $x$, External data source $D$, Generator $\mathrm{LLM}$, Retriever $R$, # of voters $m$, # of retrieval per voter $k$, Per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$, Total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$
Output: Private answer $y$
1
[…]
3 $D_{1},\ldots,D_{m}\leftarrow$ split $D$ into $m$ disjoint subset;
4
5for $t\leftarrow 1$ to $T_{\max}$ do
6       for $i\leftarrow 1$ to $m$ do
             $y_{t}^{(i)}\leftarrow\mathrm{LLM}_{t}(x,R(x,D_{i}),y_{<t})$;
              // Generate a token with RAG
7
8       end for
9      $h_{t}\leftarrow$ Build a histogram of tokens from $y_{t}^{(1)},\ldots,y_{t}^{(m)}$;
       $y_{t}\leftarrow\mathrm{LimitedDomain}(h_{t},\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$;
        // Privately choose the most frequent token
[…]
             return $(y_{1},\ldots,y_{t})$;
              // Halt if end of sequence
11
12       end if
13
14 end for
15return $(y_{1},\ldots,y_{T_{\max}})$;
16
Algorithm 1 DPVoteRAG
By the nature of retrieval in RAG—retrieving relevant documents for a question, the LLM outputs can depend on a sensitive individual’s document. Therefore, our algorithmic design needs to relax the dependency of a single individual’s document on the output, while exploiting the external data source, to achieve a reasonable privacy-utility tradeoff. Inspired by the differentially private generation via sample-and-aggregate framework in Section 2.2.2, we present a differentially private voting algorithm for RAG—DPVoteRAG.
Given a prompt $x$ and external data source $D$, DPVoteRAG first partitions $D$ into $m$ disjoint datasets $D_{1},\ldots,D_{m}$. Then, for each $i=1,\ldots,m$, it retrieves $k$ relevant documents from $D_{i}$, feeds them into the LLM along with the original prompt $x$, and generates a next token. It collects these tokens to form a histogram over the vocabulary. It remains to privately choose the most frequent element from the histogram. While it is generally hard to do so when the histogram dimension is large as in our setting, e.g., the vocabulary size of OPT [41] is 50272, there is a line of work in the community to overcome this difficulty. Following the work by [15], we integrate the LimitedDomain mechanism [8] into our algorithm. The mechanism enables us to reduce the histogram dimension significantly with some cost in privacy budget and thus achieve a better privacy-utility tradeoff. By its design, the LimitedDomain mechanism possibly outputs the designed null token. In such a case, we halt the algorithm, or equivalently, regard that it outputs the end of sequence token.  [^2] We find that by choosing the appropriate size of reduced dimension, the LimitedDomain mechanism in our experiment rarely outputs the null token. Finally, we append the chosen token to the next input to the LLM. We repeat this process until we see the end of sequence token chosen or reach the maximum number of token generation, which is computed in advance from the per-token and total privacy budget  [^3] The maximum number of token generation is computed as follows. We first calculate the maximum numbers of composition with the sequential and advanced composition theorem [12] under the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. Then, we take the maximum of two numbers of possible composition.. We present the concrete algorithm in Algorithm 1. The formal privacy analysis is as follows.
Theorem 1.
For any question $x$, DPVoteRAG satisfies $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$-DP w.r.t. the external data source $D$.
The guarantee simply follows from the privacy guarantee of the LimitedDomain mechanism and the composition theorem used to compute $T_{\max}$ in the algorithm.
3.2 DPSparseVoteRAG: Differentially Private Voting Algorithm for RAG with Sparse Vector Technique
Input: Prompt $x$, External data source $D$, Generator $\mathrm{LLM}$, Retriever $R$, # of voters $m$, # of retrieval per voter $k$, Per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$, Total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, Threshold $\tau$
Output: Private answer $y$
1
[…]
3 $D_{1},\ldots,D_{m}\leftarrow$ split $D$ into $m$ disjoint subset;
4
5$(\epsilon_{\mathrm{token-RAG}},\delta_{\mathrm{token-RAG}})\leftarrow$ $(\epsilon_{\mathrm{token}}/2,\delta_{\mathrm{token}})$;
6 $\epsilon_{\mathrm{token-Lap}}\leftarrow\epsilon_{\mathrm{token}}/2$;
7 $c\leftarrow 1$;
8 $t\leftarrow 1$;
9 $\tau_{1}\leftarrow\tau+\mathrm{Lap}(\nicefrac{{2}}{{\epsilon_{\mathrm{token-Lap}}}})$;
10 while $c\leq c_{\max}$ do
11       $y_{t}^{\mathrm{non-RAG}}\leftarrow\mathrm{LLM}_{t}(x,``^{\prime\prime},y_{<t})$;
12       for $i\leftarrow 1$ to $m$ do
             $y_{t}^{(i)}\leftarrow\mathrm{LLM}_{t}(x,R(x,D_{i}),y_{<t})$;
              // Generate a token with RAG
13
14       end for
15      $h_{t}\leftarrow$ Build a histogram of tokens from $y_{t}^{(1)},\ldots,y_{t}^{(m)}$;
16       $a_{t}\leftarrow$ Extract a count of $h_{t}$ at $y_{t}^{\mathrm{non-RAG}}$;
17       $\nu_{t}\leftarrow\mathrm{Lap}(\nicefrac{{4}}{{\epsilon_{\mathrm{token-Lap}}}})$;
18       if $a_{t}+\nu_{t}\leq\tau_{c}$ then
             /* Privately choose the most frequent token when RAG outputs disagree with non-RAG output */
19             $y_{t}\leftarrow\mathrm{LimitedDomain}(h_{t},\epsilon_{\mathrm{token-RAG}},\delta_{\mathrm{token-RAG}})$;
20             $c\leftarrow c+1$;
21             $\tau_{c}\leftarrow\tau+\mathrm{Lap}(\nicefrac{{2}}{{\epsilon_{\mathrm{token-Lap}}}})$;
22
23       end if
24      else
             /* Use non-RAG output when RAG outputs agree with it */
25             $y_{t}\leftarrow y_{t}^{\mathrm{non-RAG}}$;
26
27       end if
[…]
             return $(y_{1},\ldots,y_{t})$;
              // Halt if end of sequence
29
30       end if
31      $t\leftarrow t+1$;
32
33 end while
34return $(y_{1},\ldots,y_{t})$;
Algorithm 2 DPSparseVoteRAG
The main drawback of the aforementioned algorithm is that we need to spend a non-negligible amount of privacy budget for each token to guarantee its quality. This prevents our algorithm to generate longer answers—sometimes it can halt before it generates the actual answers due to privacy budget shortage. More concretely, consider the following question-answering example:
Question: what type of literature is the great gatsby Ground Truth Answer: novel
Here are possible outputs from (non-private) RAG and our DPVoteRAG given the retrieved documents.
RAG Output: The Great Gatsby is a novel written by American author F. Scott Fitzgerald. DPVoteRAG Output: The Great Gatsby is a
While non-private RAG correctly answers the question, due to the pre-fixed total privacy budget, DPVoteRAG can only output 5 words and thus it fails to output the ground truth answer, novel.
However, having a closer look to our voting algorithm, we observe that there is a room for improvement. When generating the 3rd word, Gatsby, every input of the LLM contains The Great, 1st and 2nd previously output words, and the great gatsby, a part of the question, even though the provided retrieved documents are different. Thus, the LLM should successfully generate Gatsby without access to the sensitive information. Ideally, we should not spend a privacy budget for such a word.
We address this by incorporating the sparse vector technique into our voting algorithm, yielding our improved algorithm DPSparseVoteRAG. In particular, before we apply private voting among generated tokens, we check if the generated tokens coincide with the token generated by the LLM without retrieved documents appended, i.e., the input is composed of the prompt and previously generated tokens only. We continue to the voting only when they do not coincide. Otherwise, we use the LLM output without retrieved documents. It is shown from the analysis of the sparse vector technique that the consumed privacy budget scales with the number of times that it uses the private voting, not with the total number of generated tokens. Consequently, the resulting algorithm, shown in Algorithm 2, enables us to spend privacy budget only when it needs sensitive information. We note that as a result of this change in our algorithm, we compute the maximum number of tokens to generate with private voting, $c_{\max}$, from the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, instead of the maximum number of token generation, $T_{\max}$, as in DPVoteRAG but in the same way to compute $T_{\max}$. The formal privacy analysis of this algorithm is as follows. The guarantee holds due to the privacy guarantee of the LimitedDomain mechanism and the NumericSparse algorithm [12].
Theorem 2.
For any question $x$, DPSparseVoteRAG satisfies $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$-DP w.r.t. the external data source $D$.
4 Experiment
We investigate how our differentially private voting RAG algorithms (Algorithms 1 and 2) work. Specifically, we ask the following questions:
  1. 1.
    How do our algorithms improve the accuracy of question-answering over non-RAG LLM while ensuring a formal privacy guarantee?
  2. 2.
    Is DPSparseVoteRAG (Algorithm 2) always a better choice than DPVoteRAG (Algorithm 1)?
  3. 3.
    Is there any useful guidance of choosing hyperparameters $m$ (the number of voters) and $\epsilon_{\mathrm{token}}$?
We study each question through extensive evaluations on the well-used benchmarking datasets with multiple LLMs.
4.1 Methodology
Datasets.
We use two question-answering benchmarking datasets for RAG: Trivia [18] and Natural Question (NQ) [21]. Each dataset consists of a list of pairs of question and answer lists, i.e., every question can have multiple answers. By following the standard evaluations in RAG [5, 23, 16], we use the Wikipedia dataset as the external data source from which a retriever finds relevant documents. For each dataset, we use a subset of 100 questions to manage the computational overhead  [^4] We first filter out questions with less than 20 ground truth documents in the Wikipedia dataset. If a question relates with only a few documents, it is inevitable that DP algorithms fail since replacing one document would change the output a lot. Then, we split remaining questions into 4 bins with 20–29, 30–39, 40–49, and 50–59 relevant documents and sample 25 questions from each bin..
Models.
The retriever we use is the Dense Passage Retriever (DPR) [20] which is built on top of BERT [6]. The DPR finds relevant documents that are close to the question query in the embedding space produced by BERT. We compare the following generator LLMs: OPT (1.3B) [41], GPT2-XL [31], and Pythia (1.4B) [3]. We additionally report the result of OPT (2.7B) in Appendix A.1.
Algorithms.
We compare our algorithms, DPVoteRAG (Algorithm 1) and DPSparseVoteRAG (Algorithm 2), with two baseline algorithms. One baseline algorithm is Non-RAG where we only provide a question to the LLM without any relevant documents appended as a prompt. In order for our algorithms to be useful, they have to outperform this baseline. The other is VoteRAG where we carry out the same voting procedure as our algorithms but choose the most frequent token across voters non-privately—the most frequent token is always chosen as the next token to generate. For each number of voters, the result of this baseline serves as the upper bound of our DP algorithms.
Experimental Setup.
We observe the results under multiple total privacy budgets, $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. More specifically, we sweep $\epsilon_{\mathrm{total}}=2$ to $40$ and set $\delta_{\mathrm{total}}=10^{-4}$. Furthermore, we consider different per-token privacy budgets for our private algorithms: $\epsilon_{\mathrm{token}}=1,2,5$ and $\delta_{\mathrm{token}}=10^{-5}$. We consider the number of voters $m$ of $10$, $20$, $30$, $40$, and $50$ for VoteRAG, and $30$, $40$, and $50$ for DPVoteRAG and DPSparseVoteRAG so as to ensure reasonable privacy-utility tradeoff and computational overhead. Too small number of voters yields too large $\epsilon_{\mathrm{token}}$ to generate accurate answers. Increasing the number of voters not only increases computational overhead but also requires a lot of relevant documents which eventually include non-relevant ones, resulting in poor utility. For DPSparseVoteRAG, we set the threshold $\tau$ to be the half of the number of voters, i.e, $\tau=m/2$. When we use the LimitedDomain mechanism to privately choose the most frequent token, we set their parameter $\bar{k}$ to be the number of voter, where $\bar{k}$ is the limited size of the domain to which we add the Gumbel noise. For voting algorithms, each voter receives 1 relevant document from corresponding subset of external knowledge dataset $D$. The utility evaluation metric is the match accuracy [25, 2, 33, 42] which measures if the prediction to a question contains any of its answers.
4.2 Results
Figure 3: Average match accuracy comparison across algorithms on Trivia (upper row) and NQ (lower row) datasets with different generator LLMs: OPT (1.3B) (left column), GPT2-XL (middle column), and Pythia (1.4B) (right column). We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$. Note that the performances of Non-RAG and VoteRAG on Trivia with Pythia (1.4B) coincide (upper right).
Comparison of algorithms.
Figure 3 shows the average match accuracy of baseline algorithms, Non-RAG and Vote-RAG, and our algorithms, DPVoteRAG and DPSparseVoteRAG for different total privacy guarantees ($\epsilon_{\mathrm{total}}$). We first observe that VoteRAG mostly improves the utility compared to Non-RAG suggesting the benefit of RAG for knowledge-extensive question-answering tasks. The only exception is the one with Pythia on Trivia—VoteRAG and Non-RAG perform comparably. We anticipate this is because Pythia is trained on Wikipedia passages and memorizes them well. Furthermore, across different datasets and generator LLMs, we observe that DPSparseVoteRAG outperforms Non-RAG mostly under $\epsilon_{\mathrm{total}}\geq 10$ when VoteRAG is better than Non-RAG. On the other hand, DPVoteRAG requires larger total privacy budget to outperform Non-RAG or can even perform worse than Non-RAG. Comparing the utility between our algorithms, we see DPSparseVoteRAG always yields better results than DPVoteRAG. This suggests the importance of the smart privacy budget usage offered by DPSparseVoteRAG.
Table 1: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.33 0.37 0.38 0.39 0.45 0.45 0.39 0.45 0.45 0.39 0.45 0.45
$\epsilon_{\mathrm{token}}=2$ 0.06 0.05 0.05 0.41 0.42 0.40 0.49 0.49 0.47 0.49 0.49 0.47
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.06 0.05 0.05 0.34 0.35 0.34 0.47 0.47 0.46
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.30 0.41 0.40 0.30 0.41 0.40 0.30 0.41 0.40 0.30 0.41 0.40
$\epsilon_{\mathrm{token}}=2$ 0.38 0.43 0.45 0.45 0.51 0.55 0.46 0.54 0.58 0.46 0.54 0.58
$\epsilon_{\mathrm{token}}=5$ 0.22 0.23 0.26 0.47 0.48 0.47 0.53 0.54 0.52 0.60 0.59 0.58
Effects of hyperparameters.
We next investigate the effects of hyperparameters of our algorithms by taking a closer look at the results in Table 1 with OPT (1.3B) on Trivia dataset under different total privacy budgets $\epsilon_{\mathrm{total}}$. We provide the detailed results with other LLMs on both Trivia and NQ in Appendix A.1.
Commonly between DPVoteRAG and DPSparseVoteRAG, we observe that the optimal $\epsilon_{\mathrm{token}}$ increases as we allow more total privacy budgets. For a tight total privacy budget, large $\epsilon_{\mathrm{token}}$ allows our algorithms to only output a few meaningful tokens; thus, smaller $\epsilon_{\mathrm{token}}$ is preferable. Conversely, under a large total privacy budget, accurate token generation with large $\epsilon_{\mathrm{token}}$ is more important than having more tokens generated with small $\epsilon_{\mathrm{token}}$. The main difference between two of our algorithms in terms of $\epsilon_{\mathrm{token}}$ is that DPSparseVoteRAG generally allows us to set larger $\epsilon_{\mathrm{token}}$ than DPVoteRAG under a fixed total privacy budget. More specifically, while setting $\epsilon_{\mathrm{token}}=2$ yields the best utility for DPVoteRAG even under $\epsilon_{\mathrm{total}}=40$, DPSparseVoteRAG performs the best with $\epsilon_{\mathrm{token}}=5$. This demonstrates the benefit of DPSparseVoteRAG to save and spend privacy budget cleverly—it can spend the saved privacy budget for generating important tokens for answering questions correctly.
With regard to the number of voters $m$ for DPVoteRAG, we generally see that having more voters results in better utility with $\epsilon_{\mathrm{token}}=1$. On the other hand, with larger $\epsilon_{\mathrm{token}}$, the number of voters has less effect on the utility and having less voters provides slightly better utility. The observation holds for all values of $\epsilon_{\mathrm{total}}$. This is due to the two distinct consequences of having more voters. More voters alleviate the effect of DP noise to the token histograms constructed within the algorithm. However, depending on the number of relevant documents to the question, there is a risk of having voters with irrelevant documents who can vote for wrong tokens. The first consequence is more dominant particularly under small $\epsilon_{\mathrm{token}}$ while the second one becomes more dominant under larger $\epsilon_{\mathrm{token}}$. For DPSparseVoteRAG, although we observe the similar trend to the above, having more voters is effective until $\epsilon_{\mathrm{token}}=2$ as opposed to $\epsilon_{\mathrm{token}}=1$ for DPVoteRAG. This is because DPSparseVoteRAG uses one half of $\epsilon_{\mathrm{token}}$ for the token generation when the RAG outputs disagree with the non-RAG output—we see the consistent impacts of the number of voters across algorithms.
4.3 Discussion
Our RAG algorithms boost the QA accuracy even under a formal privacy guarantee.
Our thorough empirical evaluation has revealed that our algorithms improves the utility of question-answering tasks over the non-RAG baseline by exploiting the external knowledge through RAG while ensuring a reasonable level of privacy $\epsilon_{\mathrm{total}}=10$. While this is mostly the case, we also observe that even (non-private) VoteRAG does not improve the utility on Trivia with Pythia. This observation implies that RAG and its private versions need to use truly sensitive data, which is completely outside the pre-training and fine-tuning data of LLMs, to be valuable enough. On the flip side, we anticipate our algorithms are even more effective for more involved tasks which heavily require the external data source for LLMs to generate meaningful responses as it is expected non-RAG LLMs perform poorly for such tasks.
DPSparseVoteRAG is strictly better than DPVoteRAG.
We further find that DPSparseVoteRAG is consistently outperforming DPVoteRAG across different LLMs and datasets. DPSparseVoteRAG augments DPVoteRAG by utilizing the non-RAG LLM and the sparse vector technique so that it only spends a privacy budget for an output token requiring sensitive external knowledge. The consistently better performances of DPSparseVoteRAG suggest the importance of separately treating token generations for meaningful tokens, i.e., tokens requiring external knowledge, and for other general tokens in the privacy-constraint setting.
$\epsilon_{\mathrm{token}}$ should allow medium-length outputs. $m$ should balance the DP noise and # of well-informed voters.
The detailed experimental results across different values of the hyperparameters give us some guidelines on how to set them properly. In particular, the per-token privacy budget, $\epsilon_{\mathrm{token}}$, should be small enough to allow the algorithms to generate a reasonably large number of tokens ($\sim 10$). As far as the resulting sequence is long enough, we should certainly make it as large as possible to enable accurate token generations. Furthermore, the number of voters, $m$, should be set by considering per-token generation quality. For a strict per-token privacy budget, e.g., $\epsilon_{\mathrm{token}}=1$, it needs to be large enough to relax the impact of the DP noise to the vote result histogram. For larger per-token privacy budgets, it should be small enough to ensure the majority of the voters receives useful external documents as the context.
5 Related Work
Privacy-preserving algorithms in large language models.
[40] proposed an empirical privacy-preserving algorithm for RAG through the synthetic data generation, while our work studies privacy-preserving RAG in the framework differential privacy, which protects the privacy of each individual documents with the theoretical guarantee. Differential privacy has been studied in many other tasks in large language models too. Prompt tuning helps tailor the LLM to new tasks from a (private) test-domain dataset. [15] and [7] study the DP mechanism on two different prompt tuning frameworks: prompt optimization and offsite prompt tuning [34]. In-context learning adapts to different tasks by illustrating some examples in the context as the task description. DP in-context learning consider the situation when the examples are picked from any private set. [35] tackles this problem by generating synthetic examples with DP and [37] solves the DP test-query by generating the answers, both in a sample-and-aggregate fashion. The differentially private pretraining and finetuning of LLMs has been studied to address the privacy concern in the training data and memory is a large bottleneck when naively deploying DP-SGD [1]. [24] focuses on the pretraining stage which introduces ghost clipping to make DP-SGD more memory efficient. [38] explores finetuning in the parameter-efficient framework LoRA [17]. Notice that DP voting plays a crucial role in these sample-and-aggregate algorithms, including ours. A basic approach is to apply the Laplacian or Gaussian mechanism [12]. [27, 28] proposed a data-dependent privacy analysis, which can be tighter when the majority vote has a large margin over other options. We integrate the LimitedDomain mechanism for our algorithm, which addresses challenges when the voting domain is large [8]; the large vocabulary size in token voting is our main bottleneck.
Composition in differential privacy.
Our algorithms generate the answers token by token, where each token needs a query to the private dataset and consumes some privacy budget. In this paper, we set up the privacy parameters before the start of the algorithm and have a pre-set maximum number of token to generate. However, the number of tokens to generate is different per question and is unknown before the algorithm starts – it is possible that the number of generated tokens is much smaller than the pre-set number but we still need to pay the full pre-defined privacy cost. A line of work [32, 14, 22, 36] tries to measure the privacy budget in fully adaptive composition where the budget consuming can interact with the data. Especially, [36] gives an analysis for this fully adaptive setting which matches the tightness of advanced composition. The idea of fully adaptive composition sounds a fit to our problem, which allows us to “pay as we go”, rather than predefining the $\varepsilon_{\max}$ before the generation process. We found the analysis for fully adaptive setting is effective for large number of steps and small budget per step, while in our algorithm the number of generated tokens would not be very large and each token generation needs a relatively large budget to guarantee the utility. This mismatch makes us stick with the advanced composition.
6 Conclusion and Future Work
We introduce the first differentially private algorithms for RAG, enabling us to enhance LLMs by domain-specific but sensitive external corpus. With our novel combination of the DP voting algorithm and sparse vector technique along with the non-private LLM, we succeed in spending privacy budget only when the LLM needs sensitive information to generate a new token. Consequently, DPSparseVoteRAG successfully generates a sufficiently long and accurate response under a reasonable privacy budget. Our experimental results demonstrate that our algorithms perform better than the non-RAG baseline across different datasets and models, showing their effectiveness.
One of our future directions is to conduct more practical empirical evaluations. The Wikipedia dataset, which we use as the external data source, is typically included in the training data of recent LLMs. RAG is particularly effective when the external knowledge is truly sensitive and thus outside the LLM training data. It is essential to conduct evaluations that are as close to the real situation. Since our usage of the sparse vector technique is applicable to any DP token generation algorithm through voting, another future direction would be to examine how it improves DP token generation across different tasks, e.g., in-context learning and prompt tuning.
Acknowledgments
TK, RW and KC would like to thank National Science Foundation NSF (CIF-2402817, CNS-1804829), SaTC-2241100, CCF-2217058, ARO-MURI (W911NF2110317), and ONR under N00014-24-1-2304 for research support.
Appendix A Additional Experimental Results
A.1 Main Results across Models and Datasets
In Figure 4, we present the average match accuracy of baseline algorithms and our algorithms for different total privacy guarantees ($\epsilon_{\mathrm{total}}$) with OPT (2.7B). We see the similar trend observed in Figure 3.
Figure 4: Average match accuracy comparison across algorithms on Trivia (left) and NQ (right) datasets with OPT (2.7B). We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$.
For completeness, we further present the detailed results, in the same form of the one provided in Table 1, with OPT (1.3B and 2.7B), GPT2-XL and Pythia (1.4B) on Trivia and NQ datasets in Tables 28.
Table 2: Average match accuracy comparison of our algorithms on NQ dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.20 0.21 0.23 0.27 0.32 0.32 0.27 0.32 0.32 0.27 0.32 0.32
$\epsilon_{\mathrm{token}}=2$ 0.03 0.03 0.03 0.26 0.23 0.23 0.36 0.35 0.36 0.36 0.35 0.36
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.03 0.03 0.03 0.24 0.22 0.22 0.33 0.31 0.30
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.11 0.20 0.28 0.11 0.21 0.28 0.11 0.21 0.28 0.11 0.21 0.28
$\epsilon_{\mathrm{token}}=2$ 0.21 0.24 0.25 0.25 0.36 0.41 0.27 0.37 0.42 0.27 0.37 0.42
$\epsilon_{\mathrm{token}}=5$ 0.07 0.07 0.07 0.26 0.25 0.26 0.40 0.35 0.37 0.46 0.43 0.47
Table 3: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.34 0.38 0.38 0.44 0.48 0.51 0.44 0.48 0.51 0.44 0.48 0.51
$\epsilon_{\mathrm{token}}=2$ 0.03 0.03 0.03 0.39 0.40 0.40 0.50 0.52 0.54 0.50 0.52 0.54
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.03 0.03 0.03 0.33 0.30 0.28 0.48 0.48 0.48
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.36 0.48 0.57 0.36 0.48 0.57 0.36 0.48 0.57 0.36 0.48 0.57
$\epsilon_{\mathrm{token}}=2$ 0.45 0.51 0.52 0.58 0.63 0.68 0.60 0.64 0.70 0.60 0.64 0.70
$\epsilon_{\mathrm{token}}=5$ 0.28 0.34 0.38 0.49 0.54 0.55 0.60 0.65 0.66 0.63 0.69 0.71
Table 4: Average match accuracy comparison of our algorithms on NQ dataset with OPT (2.7B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.24 0.29 0.32 0.34 0.42 0.45 0.34 0.42 0.45 0.34 0.42 0.45
$\epsilon_{\mathrm{token}}=2$ 0.03 0.03 0.03 0.32 0.32 0.32 0.44 0.47 0.46 0.44 0.47 0.46
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.03 0.03 0.03 0.30 0.30 0.30 0.37 0.36 0.37
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.15 0.18 0.24 0.15 0.19 0.25 0.15 0.19 0.25 0.15 0.19 0.25
$\epsilon_{\mathrm{token}}=2$ 0.25 0.29 0.33 0.36 0.46 0.52 0.37 0.48 0.54 0.37 0.48 0.54
$\epsilon_{\mathrm{token}}=5$ 0.11 0.09 0.10 0.33 0.35 0.34 0.51 0.52 0.50 0.53 0.56 0.57
Table 5: Average match accuracy comparison of our algorithms on Trivia dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.29 0.33 0.34 0.32 0.39 0.42 0.32 0.39 0.42 0.32 0.39 0.42
$\epsilon_{\mathrm{token}}=2$ 0.16 0.16 0.16 0.37 0.36 0.35 0.45 0.44 0.43 0.45 0.44 0.43
$\epsilon_{\mathrm{token}}=5$ 0.03 0.03 0.03 0.17 0.16 0.17 0.34 0.31 0.32 0.43 0.41 0.42
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.32 0.35 0.37 0.32 0.35 0.37 0.32 0.35 0.37 0.32 0.35 0.37
$\epsilon_{\mathrm{token}}=2$ 0.37 0.34 0.42 0.42 0.44 0.54 0.43 0.46 0.56 0.43 0.46 0.56
$\epsilon_{\mathrm{token}}=5$ 0.24 0.25 0.24 0.41 0.41 0.42 0.54 0.51 0.52 0.57 0.56 0.58
Table 6: Average match accuracy comparison of our algorithms on NQ dataset with GPT2-XL under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.20 0.30 0.30 0.23 0.33 0.34 0.23 0.33 0.34 0.23 0.33 0.34
$\epsilon_{\mathrm{token}}=2$ 0.14 0.14 0.15 0.33 0.33 0.32 0.37 0.37 0.36 0.37 0.37 0.36
$\epsilon_{\mathrm{token}}=5$ 0.04 0.04 0.05 0.13 0.14 0.16 0.29 0.30 0.31 0.35 0.37 0.36
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.09 0.19 0.19 0.09 0.19 0.20 0.09 0.19 0.20 0.09 0.19 0.20
$\epsilon_{\mathrm{token}}=2$ 0.24 0.26 0.32 0.27 0.34 0.36 0.28 0.36 0.38 0.28 0.36 0.38
$\epsilon_{\mathrm{token}}=5$ 0.11 0.12 0.09 0.32 0.31 0.32 0.39 0.41 0.39 0.42 0.43 0.42
Table 7: Average match accuracy comparison of our algorithms on Trivia dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.23 0.26 0.27 0.34 0.39 0.41 0.34 0.39 0.41 0.34 0.39 0.41
$\epsilon_{\mathrm{token}}=2$ 0.14 0.14 0.14 0.29 0.30 0.30 0.43 0.46 0.44 0.43 0.46 0.44
$\epsilon_{\mathrm{token}}=5$ 0.03 0.03 0.03 0.14 0.14 0.14 0.27 0.27 0.27 0.38 0.39 0.38
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.26 0.31 0.39 0.26 0.31 0.39 0.26 0.31 0.39 0.26 0.31 0.39
$\epsilon_{\mathrm{token}}=2$ 0.38 0.37 0.40 0.45 0.47 0.48 0.46 0.52 0.53 0.46 0.52 0.53
$\epsilon_{\mathrm{token}}=5$ 0.22 0.23 0.24 0.40 0.39 0.38 0.52 0.52 0.51 0.56 0.57 0.56
Table 8: Average match accuracy comparison of our algorithms on NQ dataset with Pythia (1.4B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
Algorithm $\epsilon_{\mathrm{total}}=5$ $\epsilon_{\mathrm{total}}=10$ $\epsilon_{\mathrm{total}}=20$ $\epsilon_{\mathrm{total}}=40$
$m$ 30 40 50 30 40 50 30 40 50 30 40 50
DPVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.16 0.19 0.19 0.20 0.24 0.25 0.20 0.24 0.25 0.20 0.24 0.25
$\epsilon_{\mathrm{token}}=2$ 0.11 0.10 0.10 0.24 0.24 0.25 0.30 0.30 0.30 0.30 0.30 0.30
$\epsilon_{\mathrm{token}}=5$ 0.02 0.02 0.02 0.11 0.11 0.11 0.22 0.22 0.22 0.30 0.31 0.31
DPSparseVoteRAG
$\epsilon_{\mathrm{token}}=1$ 0.04 0.12 0.16 0.04 0.12 0.18 0.04 0.12 0.18 0.04 0.12 0.18
$\epsilon_{\mathrm{token}}=2$ 0.18 0.19 0.19 0.27 0.34 0.33 0.31 0.38 0.39 0.31 0.38 0.39
$\epsilon_{\mathrm{token}}=5$ 0.05 0.06 0.06 0.23 0.22 0.22 0.36 0.40 0.37 0.45 0.47 0.45
A.2 Further Results on Trivia Dataset with OPT (1.3B)
Effects of Number of Ground Truth Relevant Documents.
Figure 5: Average match accuracy for questions with different numbers of ground truth relevant documents across baseline algorithms and our algorithms with OPT (1.3B) on Trivia dataset.
Figure 5 shows the performances for different numbers of ground truth relevant documents. We see questions with more relevant documents tend to be answered correctly across algorithms, even with Non-RAG. We also observe that the benefit of having more voters (larger $m$) is larger for small $\epsilon_{\mathrm{token}}$. Interestingly, the benefit is significant when there are less relevant documents for DPVoteRAG and when there are more relevant documents for DPSparseVoteRAG. We don’t see significant difference between $m=30$ and $50$ when $\epsilon_{\mathrm{token}}=5$.
Number of Generated Tokens.
Figure 6: Average numbers of generated tokens for each $\epsilon_{\mathrm{token}}$ and $\epsilon_{\mathrm{total}}$ with OPT (1.3B) on Trivia dataset. We fix $m=50$.
Figure 6 shows the numbers of tokens generated by DPVoteRAG and DPSparseVoteRAG. As we expect by the design of DPSparseVoteRAG, we see DPSparseVoteRAG generates much more tokens than DPVoteRAG. This implies the effectiveness of the sparse vector technique in DPSparseVoteRAG to smartly spend privacy budget enabling long enough token sequences.
Voting Frequency of DPSparseVoteRAG.
Figure 7: Histograms of voting frequency of DPSparseVoteRAG for each $\epsilon_{\mathrm{token}}$ with OPT (1.3B) on Trivia dataset. We fix $\epsilon_{\mathrm{total}}=20$ and $m=50$.
Figure 7 demonstrates how often DPSparseVoteRAG carries out the voting procedure. We see that DPSparseVoteRAG uses the voting procedure only for the 10% of tokens in the total output for most of the questions when $\epsilon_{\mathrm{token}}\geq 2$. On the other hand, when $\epsilon_{\mathrm{token}}=1$, it uses the voting more often. The results validate that the sparse vector technique helps spend a privacy budget in a principled manner. Furthermore, the results imply that small $\epsilon_{\mathrm{token}}$ can lead to the unnecessary use of the voting, degrading the overall performance.